# A Hardware-Friendly Shuffling Countermeasure Against Side-Channel Attacks for Kyber

### v1

@@ Abstract
CRYSTALS-Kyber (a.k.a. Kyber) has been drafted to be standardized as the only key encapsulation mechanism (KEM) scheme by the national institute of standards and technology (NIST) to withstand attacks by large-scale quantum computers. However, the side-channel attack (SCA) on its implementation is still needed to be well considered for the upcoming migration. In this brief, we propose a secure and efficient hardware implementation for Kyber by incorporating a novel compact shuffling architecture. First of all, we modify the Fisher-Yates shuffle to make it more hardware-friendly. We then design an optimized shuffling architecture for the well-known open-source Kyber hardware implementation to enhance the security of all the potential side-channel leakage points. Finally, we implement the modified Kyber design on FPGA and evaluate its security and performance. The security is verified by conducting the correlation power analysis (CPA) attacks on the hardware. Meanwhile, FPGA place-and-route results show that the proposed design reports only 8.7% degradation on the hardware efficiency compared with the original unprotected version, much better than existing hiding schemes.
@@ I Introduction
With the rapid development of quantum computing, modern cryptography algorithms are facing serious threats [1]. Since 2016, the national institute of standards and technology (NIST) has launched a standardization competition for post-quantum cryptography (PQC) and solicited proposals from around the world [2], aiming to form new PQC standards ahead of the coming quantum era. During the PQC standardization process, the lattice-based cryptography (LBC) becomes the most competitive category, occupying three of all four winners to be standardized after three rounds of evaluation [3]. CRYSTALS-Kyber (Kyber) is the only key encapsulation mechanism (KEM) protocol among them. During the upcoming transition of existing conventional cryptographic algorithms to Kyber, it is essential to evaluate the security against physical attacks, such as the side-channel attack (SCA).
In recent years, researchers have proven that both software and hardware implementations of Kyber have the possibility for side-channel leakage, and have successfully implemented SCAs [4, 5, 6]. On the software side, Bock et al[4] successfully recovered the key from the polynomial multiplication in Kyber’s decryption procedure, and Sim et al[5] proved that the disclosure of sensitive variables in Barrett reduction would have exposed the entire key. On the hardware side, Zhao et al[6] recently pointed out that there are three side-channel leakage points in Kyber’s decryption procedure, and successfully completed SCA on the open-source hardware implementation from [7].
@@
However, defense work for Kyber, especially on hardware platform, is still far from sufficient, where masking and hiding both are usually adopted countermeasures. Zhao et al[6] and Kamucheka et al[8] proposed two different hardware masking schemes, both of which introduce significant area overhead. On the other hand, two hiding schemes on hardware platform are provided in [9] and [10]. In [9], Morait et al. used the method of randomized clock and duplication to reduce the energy correlation during the operation of Kyber. In [10], Jati et al. adopted the methodology of adding random delays and address & instruction shuffling to reduce the energy correlation. Nevertheless, these two hiding schemes result in high resource overhead and clock cycle overhead, respectively. In this brief, we try to propose a hiding protection method for Kyber’s hardware implementation by taking the performance and security both into consideration.
@@
This work provides a good tradeoff between security enhancement and hardware efficiency. Specifically, we modify the Fisher-Yates shuffle [11] to make it more hardware-friendly. Based on the algorithm, we propose an optimized shuffling architecture and apply it to the open-source hardware implementation of Kyber from [7]. We shuffle the sub-operation orders of all the potential side-channel leakage points during Kyber’s decryption procedure. The experimental results show that the proposed design can effectively improve the side-channel security, and reports only 8.7% degradation on the hardware efficiency compared with the original unprotected version.
@@ II-A Related Work on Shuffling
As an efficient and effective countermeasure, shuffling is widely used in various cryptographic implementations against SCAs. In recent years, numerous works have been conducted on the security research pertaining to shuffling in the software implementations of lattice-based cryptography. In contrast, there are relatively few studies on the use of shuffling on hardware implementations, as evidenced by the limited entries in Table I. In [10], Jati et al. adopted address & instruction shuffling to reduce the energy correlation. In [12], Chen et al. proposed a low-cost shuffling countermeasure against side-channel attacks for NTT-based post-quantum cryptography. The random permutations used in these two schemes are generated outside of the hardware and preloaded into storage in advance. These schemes require lots of storage space to achieve high security. For example, storing $1\times 10^{8}$ permutations of length 64 would require $3.8\times 10^{10}$ bits of storage space. In [13], Bayrak et al. proposed a shuffling architecture to generate random permutations. However, under the premise of generating the same size of random permutations, the area consumed by their architecture is almost 4x that of the proposed architecture (only the random permutation generator).
TABLE I: Comparison with Existing Shuffling Methods
Work Full Permutation No Dynamic Input Required No Pre-Storage Required
@@
TACO [13] $\checkmark$ $\times$ $\checkmark$
@@
II-B Side-Channel Leakage Points in Kyber
The whole Kyber protocol consists of three procedures, which are key generation, encryption, and decryption. In a cryptographic protocol, the operations that involve both the sensitive values and the intermediate values of plaintext/ciphertext are more prone to side-channel leakage [14]. In the Kyber protocol, the decryption procedure contains multiple operations that satisfy this condition. The main operations of Kyber’s decryption procedure are as follows:
@@ II-B Side-Channel Leakage Points in Kyber
The operations in (1) include number theoretic transform (NTT), point-wise multiplication (PWM), inverse number theoretic transform (INTT), subtraction, and compress, where $m$ denotes the decrypted message, $\boldsymbol{\hat{s}}$ denotes the unpacked secret key, $\boldsymbol{u}$ and $v$ denote the unpacked ciphertext. NTT, PWM, and INTT are used to reduce the complexity of polynomial multiplication. In Kyber’s decryption procedure, Compressq is a function that compresses a polynomial coefficient to 0 or 1, where $q$ stands for the modulo $3329$. According to [6], there are three side-channel leakage points in Kyber’s decryption procedure (hardware) as shown in (2).
@@
The three leakage points lie in PWM, the modular reduction after PWM, and the subtraction after INTT, respectively. In addition, INTT has been successfully attacked in software implementations [15], so we retain the protection of INTT.
@@ III Proposed Shuffling Architecture for Kyber
Fisher-Yates shuffle [11] is one of the most widely used shuffling algorithms. When it is implemented in hardware, two problems need to be solved. One is to continuously input random seeds and the other is to dynamically reduce the range of indexes. In this work, we have elaborately solved these problems. The key ideas of the improved algorithm will be shown in Algorithm 1.
@@ III-A Random Permutation Generator
Random permutation generator (RPG) is a module that generates random permutations from an ordered one. There are five different random permutations to be generated, which are $00\sim 3f$, $40\sim 7f$, $00\sim 7f$, $80\sim bf$, and $c0\sim ff$. Since there is a correspondence between these permutations, it is only necessary to generate the permutation $00\sim 3f$, and the other permutations can be obtained by extending it.
@@
Note that the first 11 elements from the left of this permutation are designed to satisfy the restrictions shown in Fig. 1(c). Actually, the designed address permutation only needs to restrict those 11 elements and can be in any form. For simplicity, we define it in the above form in this brief.
@@
where the value 28 is in the hexadecimal format. The variable rest denotes the number of the remaining elements to be selected minus 1, which is computed by a decrement counter. The output signal $idx^{\prime}$ is selected from $idx\_0$, $idx\_1$, and $rest$, served as the control signal of MUX0 and DEMUX in (a). Part (d) contains a LFSR (used to continuously generate random numbers) and a buffer. The buffer outputs every six cycles to convert six 1-bit numbers into a 6-bit number. When $rst\_l$ is equal to 1, LFSR updates the registers by using the input random seed. Part (e) contains a controller made up with several logic circuits. It produce the control signals $ena$, $sel\_0$, $sel\_1$, and $finish$ based on the input signals $rst$ and $start$ for the other three modules.
@@
The process of RPG will not increase the consumption of total time of Kyber with this protection since RPG is conducted in parallel with Kyber, costs much less cycles, and has shorter critical path. So we design it in a compact way to simplify the control and save area.
@@ III-B Address Controller
Figure 4: The CPA attacks results of the unprotected hardware implementation (left) and the protected hardware implementation (right).
@@ IV-A Evaluation on SCA Resistance
To better assess the level of improvement in side-channel security after applying the protective measures, we conduct correlation power analysis (CPA) attacks (one of the most common SCAs) [16] on PWM for both the unprotected and protected hardware implementations, and the experimental results are shown in Fig. 4. The experimental results indicate that the unprotected Kyber’s hardware implementation can be successfully attacked with a minimum of $0.9k$ power traces, while the protected Kyber’s hardware implementation does not show significant correlation even with $100k$ power traces. It can be concluded that the attack complexity is increased by more than 100x for the attacker. It should be pointed out that the rest leakage points have the similar conclusions as they adopted the same defense countermeasure. We have omitted the experimental results of them here for simplicity.
TABLE II: Comparison of the FPGA Area and Time Performance
Hardware Implementation Parameter FPGA Platform LUTs FFs Slices DSPs BRAMs ENS 1 Cycles (×$\times 10^{3}$) Frequency (MHz) ATP 2 (ENS×$\times ms$)
Xing et al[7]3 (unprotected) Kyber512 Artix-7 7,353 4,633 2,173 2 3 2,973 6.7 206 96.7
Kamucheka et al[8] Kyber512 Virtex-7 143,112 - 81,746 60 294 146,546 126.6 100 186×$\times 10^{3}$
Jati et al[10] Kyber512 Artix-7 7,151 3,730 2,260 4 2 5.5 3,560 57.2 258 789.3
This work Kyber512 Artix-7 8,143 5,151 2,433 2 3 3,233 6.7 206 105.2
Xing et al[7]3 (unprotected) Kyber768 Artix-7 7,353 4,633 2,173 2 3 2,973 10.0 206 144.3
Moraitis et al[9] Kyber768 Artix-7 14,341 9,190 4,734 4 $\geq$2 6 $\geq$6,134 10.0 $\leq$206 $\geq$297.8
This work Kyber768 Artix-7 8,143 5,151 2,433 2 3 3,233 10.0 206 156.9
  • 1. ENS (equivalent number of slices) = Slices + DSPs $\times$ 100 + BRAMs $\times$ 200 [17].
  • 3. We reevaluate this project using Artix-7 XC7A100T for a fair comparison.
  • 4. The number of Slices is approximately computed by using LUTs/4 + FFs/8 [17].
@@ IV-B Comparisons on Area and Performance
Table II shows the experiment results and comparison of Kyber’s hardware implementations with different hiding schemes, including the area and time performance. In order to take the DSPs and BRAMs into consideration, we use the equivalent number of slices (ENS) as a new metric for area with the same approximation method used in [17]. For the hardware efficiency, the ATP is computed as ENS $\times$ Time and listed in the rightmost column of the table. Compared to the unprotected Kyber implementation [7], the proposed protected design only consumes an extra resource of 260 Slices. The numbers of DSPs and BRAMs both are the same. Compared with the unprotected design, the ATP (also area) of the proposed version is increased by only 8.7%.
@@ V Conclusion
In this brief, we have devised an optimized shuffling architecture and integrated it into the Kyber’s hardware implementation to achieve shuffling of all the potential side-channel leakage points during Kyber’s decryption procedure. The experimental results show that the proposed design can effectively improve the side-channel security, and reports only 8.7% degradation on the hardware efficiency compared with the original unprotected version, much better than existing hiding schemes.

### v2

@@ Abstract
CRYSTALS-Kyber has been standardized as the only key-encapsulation mechanism (KEM) scheme by NIST to withstand attacks by large-scale quantum computers. However, the side-channel attacks (SCAs) on its implementation are still needed to be well considered for the upcoming migration. In this brief, we propose a secure and efficient hardware implementation for Kyber by incorporating a novel compact shuffling architecture. First of all, we modify the Fisher-Yates shuffle to make it more hardware-friendly. We then design an optimized shuffling architecture for the well-known open-source Kyber hardware implementation to enhance the security of all known and potential side-channel leakage points. Finally, we implement the modified Kyber design on FPGA and evaluate its security and performance. The security is verified by conducting correlation power analysis (CPA) and test vector leakage assessment (TVLA) on the hardware. Meanwhile, FPGA place-and-route results show that the proposed design reports only 8.7% degradation on the hardware efficiency compared with the original unprotected version, much better than existing hardware hiding schemes.
@@ I Introduction
With the rapid development of quantum computing, modern cryptography faces severe threats. Since 2016, NIST has held a post-quantum cryptography (PQC) standardization competition, receiving global proposals. After three rounds, CRYSTALS-Kyber (Kyber) was chosen as the only key-encapsulation mechanism (KEM) protocol [1]. Recently, NIST issued FIPS 203, a standard for module-lattice-based key-encapsulation mechanism (ML-KEM) based on CRYSTALS-Kyber. As of now, research on the efficient hardware implementations of ML-KEM/Kyber has attained a relatively mature stage [2, 3]. Nevertheless, during the impending shift from existing conventional cryptographic algorithms to ML-KEM/Kyber, it is equally crucial to assess their security against physical attacks, particularly side-channel attacks (SCAs) and fault injection attacks (FIAs).
SCAs recover keys by acquiring power consumption or electromagnetic emanations from cryptographic devices, and typically employ countermeasures such as hiding or masking to defend against them. FIAs, on the other hand, aim to recover keys by introducing faults into cryptographic devices during their operation. Error detection mechanisms can effectively identify and respond to the injection of these faults. On the SCAs side, Tosun et al[4] successfully recovered the full key of Kyber from the polynomial multiplication on Kyber’s software implementations. Zhao et al[5] recently pointed out that there are three side-channel leakage points in Kyber’s decryption procedure, and successfully completed SCAs on Kyber’s hardware implementations. On the FIAs side, Ni et al[6] successfully implemented bitstream fault injection attacks on Kyber’s hardware implementations. In addition, combined fault and power attacks also pose a significant threat to lattice-based cryptographic algorithms.
@@
However, defense work for Kyber, especially on hardware platform, is still far from sufficient. Sarker et al[7] proposed the error detection architectures for hardware/coftware co-design approaches of number theoretic transform (NTT). Zhao et al[5] and Kamucheka et al[8] proposed two different hardware masking schemes, both of which introduce significant area overhead. On the other hand, two hiding schemes on hardware platform are provided in [9] and [10]. In [9], Morait et al. used the method of randomized clock and duplication to reduce the energy correlation during the operation of Kyber. In [10], Jati et al. adopted the methodology of adding random delays and address & instruction shuffling to reduce the energy correlation. Nevertheless, these two hiding schemes result in high resource overhead and clock cycle overhead, respectively. In this brief, we try to propose a hiding protection method against SCAs for Kyber’s hardware implementation by taking the performance and security both into consideration.
@@
This work provides a good tradeoff between security enhancement and hardware efficiency. Specifically, we modify the Fisher-Yates shuffle to make it more hardware-friendly. Based on the algorithm, we propose an optimized shuffling architecture and apply it to the open-source hardware implementation of Kyber from [11]. We shuffle the sub-operation orders of all the potential side-channel leakage points during Kyber’s decryption procedure. The proposed design shows only an 8.7% efficiency drop and surpasses current hiding schemes.
@@ II-A Related Work on Shuffling
As an efficient and effective countermeasure, shuffling is widely used in various cryptographic implementations against SCAs. In [12], Ravi et al. proposed three variants of the shuffling countermeasure with varying granularity for the NTT on software platform. In [10, 13, 14], three different shuffling countermeasures on the hardware platform are proposed, as shown in Table I. The schemes from [10, 13] require lots of storage space to achieve high security as their random permutations are generated outside of the hardware and preloaded into storage in advance. Therefore, they require lots of storage space to achieve high security. Differently, Zijlstra et al[14] used the random permutation generator (RPG) from [15] to protect the lattice-based cryptographic schemes. However, the RPG they proposed is quite costly, making it difficult to meet the low-cost requirements.
TABLE I: Comparison with Existing Hardware Shuffling Methods
Work Huge Space of Permutations No Dynamic Input Required No Pre-Storage Required
@@
INDOCRYPT [14] $\checkmark$ $\times$ $\checkmark$
@@
II-B Kyber and Side-Channel Leakage Points in Kyber
Kyber is a CCA-secure protocol whose security is based on the module learning with errors (MLWE) problem. The core operations of Kyber are polynomial multiplications over the polynomial ring $\mathcal{R}_{q}=\mathbb{Z}_{q}/(X^{n}+1)$, where $n=256$, $q=3329$, and $k=2$, $3$, and $4$. The whole Kyber protocol consists of three procedures, which are key generation, encryption, and decryption. The side-channel leakage points of the secret key are mainly concentrated in Kyber’s decrypting procedure, and the main operations of this procedure are as follows:
@@ II-B Kyber and Side-Channel Leakage Points in Kyber
They include NTT, point-wise multiplication (PWM), inverse NTT (INTT), subtraction, and compress, where $m$ denotes the decrypted message, $\boldsymbol{\hat{s}}$ denotes the unpacked secret key, $\boldsymbol{u}$ and $v$ denote the unpacked ciphertext. NTT, PWM, and INTT are used to reduce the complexity of polynomial multiplications from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\cdot logn)$.
According to [5], there are three side-channel leakage points in Kyber’s decryption procedure (hardware) as shown in (2).
@@
The three leakage points lie in PWM, the modular reduction after PWM, and the subtraction after INTT, respectively. In addition, INTT has been successfully attacked in software implementations [16], so we retain the protection of INTT by shuffling the butterfly units using a series of fixed-length random permutations produced by the proposed RPG.
@@ III Proposed Shuffling Architecture for Kyber
Fisher-Yates shuffle is one of the most widely used shuffling algorithms. When it is implemented in hardware, two problems need to be solved. One is to continuously input random seeds and the other is to dynamically reduce the range of indexes. In this work, we have elaborately solved these problems. The key ideas of the improved algorithm will be shown in Algorithm 1.
@@ III-A Random Permutation Generator
RPG is a module that generates random permutations from an ordered one. There are five different random permutations to be generated, which are $00\sim 3f$, $40\sim 7f$, $00\sim 7f$, $80\sim bf$, and $c0\sim ff$. Since there is a correspondence between these permutations, it is only necessary to generate the permutation $00\sim 3f$, and the other permutations can be obtained by extending it.
@@
Actually, the designed address permutation only needs to restrict those 11 elements and can be in any form. For simplicity, we define it in the above form in this brief.
@@
where the value 28 is in the hexadecimal format. The variable rest denotes the number of the remaining elements to be selected minus 1, which is computed by a decrement counter. The output signal $idx^{\prime}$ is selected from $idx\_0$, $idx\_1$, and $rest$, served as the control signal of MUX0 and DEMUX in (a). Part (d) contains a linear feedback shift register (LFSR) with depth of 32 and a buffer. The buffer outputs every six cycles to convert six 1-bit numbers into a 6-bit number (index). The random seeds are generated by an external true random number generator (TRNG). When $rst\_l$ is equal to 1, the LFSR will be updated by the input random seed. Part (e) contains a controller made up with several logic circuits. It produce the control signals $ena$, $sel\_0$, $sel\_1$, and $finish$ based on the input signals for the other three modules.
@@
The protection will not increase the consumption of total time of Kyber since RPG is conducted in parallel with Kyber, costs much less cycles, and has shorter critical path.
@@
TABLE II: Comparison of the FPGA Area and Time Performance
Implementation Parameter Platform LUTs FFs Slices DSPs BRAMs ENS 1 Cycles (×$\times 10^{3}$) Frequency (MHz) ATP 2 (ENS×$\times ms$)
Xing et al[11]3 Kyber512 Artix-7 7353 4633 2173 2 3 2973 6.7 206 96.7
Kamucheka et al[8] Kyber512 Virtex-7 143112 - 81746 60 294 146546 126.6 100 186×$\times 10^{3}$
Jati et al[10] Kyber512 Artix-7 7151 3730 2260 4 2 5.5 3560 57.2 258 789.3
This work Kyber512 Artix-7 8143 5151 2433 2 3 3233 6.7 206 105.2
Xing et al[11]3 Kyber768 Artix-7 7353 4633 2173 2 3 2973 10.0 206 144.3
Moraitis et al[9] Kyber768 Artix-7 14341 9190 4734 4 $\geq$2 6 $\geq$6134 10.0 $\leq$206 $\geq$297.8
This work Kyber768 Artix-7 8143 5151 2433 2 3 3233 10.0 206 156.9
  • 1. ENS (equivalent number of slices) = #Slices + 100 $\times$ #DSPs + 200 $\times$ #BRAMs  [17].
  • 3. The unprotected work was originally evaluated using Artix-7 XC7A12T. We reevaluate it using Artix-7 XC7A100T for a fair comparison.
  • 4. The number of Slices is approximately computed by using 0.25 $\times$ #LUTs + 0.125 $\times$ #FFs [17].
Figure 4: The CPA results of the unprotected scheme with $4\times 10^{3}$ EM traces (a) and the protected scheme with $1\times 10^{5}$ EM traces (b). The TVLA results of the unprotected scheme with $1\times 10^{4}$ EM traces (c) and the protected scheme with $1\times 10^{7}$ EM traces (d). The red dashed lines represent ±$\pm 4.5$.
@@ IV-A Evaluation on SCA Resistance
To better assess the level of improvement in side-channel security after applying the protective measures, we conduct correlation power analysis (CPA) on PWM and test vector leakage assessment (TVLA) on Kyber’s decryption procedure. In the experiments, Pearson correlation coefficient and Welch’s t-test are used as metrics in CPA and TVLA, respectively. We deploy our experiments for Kyber768 on a target board containing an Artix-7 XC7A100T FPGA.
We target the registers at the output ports of multipliers and utilize the Hamming distance model to simulate the energy consumption resulting from the dynamic flipping of these registers, similar to the methodology employed in [5]. Finally, the correlation between the assumed energy values and the collected energy values are calculated to analyze the correct keys. As shown in Fig. 4, all experiments have achieved the expected results, i.e., the proposed shuffling countermeasure with huge permutation space can significantly improve the side-channel security of Kyber’s hardware implementation. It should be pointed out that we have omitted the CPA results of modular reduction and subtraction here for simplicity as they adopt the same defense countermeasure and have the similar conclusions.
@@ IV-B Comparisons on Area and Performance
Table II shows the experiment results and comparison of Kyber’s hardware implementations with different hiding schemes, including the area and time performance. Compared to the unprotected Kyber implementation [11], the proposed protected design only consumes an extra resource of 260 Slices. The numbers of DSPs and BRAMs both are the same. Compared with the unprotected design, the ATP (also area) of the proposed version is increased by only 8.7%.
@@ V Conclusion
In this brief, we have devised an optimized shuffling architecture against SCAs for Kyber’s hardware implementation. The experimental results show that the proposed design can effectively improve the side-channel security, and reports only 8.7% degradation on the hardware efficiency compared with the original unprotected version, much better than existing hiding schemes.
VI Discussion
For lattice-based KEMs and signature algorithms, NTT-based polynomial multiplications serve as their most basic operation, with similar security foundations, making the proposed method applicable to other lattice-based schemes. In addition, guarding against FIAs is equally important as guarding against SCAs. However, this brief has so far only protected the Kyber algorithm and has not yet integrated error detection schemes. Future work will apply the proposed method to other promising algorithms like Dilithium and Raccoon, and consider combining it with error detection schemes.